The current U.S. dollar-yen spot rate is 125.00\%/\$ . If the 90-day forward exchange rate is 127.60\%/\$ then the yen is selling at a per annum ___of_
Answer #1

The formula to calculate forward premium is given below:

Forward premium = Forward rate -Spot rate Forward rate 360 Number of days - 1x100

Substitute 127.60 for forward rate, 125 for spot rate, 90 for number of days,

(127.60–125 x 360 x100 Forward premium = 127.60 *100*100 2.60 -> 4x100 127.60 =0.020376x4x100 = 8.15%

Hence, the yen is selling at 8.15% discount per annum.